OpenStudy (anonymous):

Your office is 60 km from your home .you drive the first 30 km at a constant 20 m/sec and the next 30 km at a constant 40.0 m/sec. what is your average speed for the entire speed for entire journey

OpenStudy (anonymous):

t=d/r\[t\text{=}\frac{30000}{20}+\frac{30000}{40}\]t=2250 r t = d r 2250 = 60000\[r = 26 \frac{2}{3} \text{ meters}/\text{second} \]